OpenStudy (anonymous):

Genes B-blonde b-brunette Mon=Bb Dad-Bb if they had 50 children,how many would have brunette hair?

OpenStudy (anonymous):

In order to solve this problem, we would need to make a Punnett Square. Here already have the ALLELES -- (B) and (b). ALLELES are letters, that stand for a trait. In this case, CAPITAL B stands for BLONDE. And LOWERCASE b stands for BRUNETTE. We also already have the GENOTYPE of each of the two parents. MOM's genotype is (B b). DAD's genotyps is also (B b). A GENOTYPE is a group of letters. A group of alleles, you see? We need to make a Punnett Square, in order to cross these genotypes. Let's put mom on TOP, and dad on the LEFT SIDE. Here's what our Punnett Square would look like, after it's been completed:  So, again, you can see that MOM is on TOP, and DAD is on the LEFT. And now, we have four squares, filled in with four different genotypes. ~~~~~~~~~~~~~~~ Before we go on, I just wanna explain how to read a genotype. In a genotype, the CAPITAL letter always matters more than the lowercase letter. For example, look at this genotype: (B b). A person with that genotype would be BLONDE, because Capital B stands for blonde, you see? It doesn't matter that the lowercase b is there, because the Capital letter always matters more. IF there is no Capital letter, THEN the lowercase letter would matter, like in (b b). ~~~~~~~~~~~~~~~~~~ Okay, so again, we've made our Punnett Square, and we've filled in all four of the blank squares. We now have four different genotypes. What that means is that Mom and Dad's children could have ANY ONE of those four genotypes. Here's a big, BIG fact: **Punnett Squares are used to predict PROBABILITY.** ~~~~~~~~~~~~~~ QUESTION: What is the probability that the child will have the genotype (B B)? ANSWER: Since there are 4 squares in all, and only 1 of the squares is (B B), the probability that the child will be (B B) is 25%. That's because 1/4 = 25% You get it? ~~~~~~~~~~~~~~ Another QUESTION: What's the probability that the child will have BLONDE hair? ANSWER: There are THREE genotypes, on the Punnett Square, for blonde hair: (B B) (B b) and another (B b) And there are FOUR squares in all. 3/4 = 75% You see? ~~~~~~~~~~~~~~ Now, let's look at your original question: "If they had 50 children,how many would have BRUNETTE hair?" Ohhh, so we're looking at BRUNETTE genotype now. There's only one genotype, on the Punnett Square, that would give a child BRUNETTE hair: (b b). So! What's 1/4? That would be 25%. And the question says that there will be 50 children. So what's 25% of 50? We can take use our calculator: 0.25×50=12.5 Our answer is 12.5 children. 12 and a half children would have BRUNETTE hair. And since that'd be weird, we can round it up to 13. 13 children. You get it?